# TCLC-GS: Tightly Coupled LiDAR-Camera Gaussian Splatting for Surrounding Autonomous Driving Scenes

### v1

@@ 11 Video Demo
We attached a Video Demo 1 and Video Demo 2 to illustrate the rendered images and depths by our TCLC-GS, alongside the ground truth (GT) data, in two different urban scenes. In these videos, the first row displays the images rendered by TCLC-GS, the second row shows the GT images, the third row presents the dense depth rendered by TCLC-GS, and the fourth row illustrates the GT sparse depth obtained by projecting LiDAR data onto the surrounding images.
We attached a Video Demo 3 to showcase a more challenging rendering scenario. This video presents image and depth rendering from a new generated ego-car trajectory which is 0.5 meter higher or 0.5 meter lower than the original ego-car trajectory. The first and third rows display the rendered RGB and depth images from a trajectory 0.5m lower than the original, while the second and fourth rows show the rendered RGB and depth images from a trajectory 0.5m higher than the original. This demonstration highlights our method’s potential to address the domain gap in data reuse caused by different hardware car settings, such as collecting data from a sedan but deploying it on an SUV.

### v2

@@ 4 Experiments
Ablation Study: To demonstrate the individual effectiveness of each component in our method, we conducted ablation studies using the Waymo dataset. Since our contributions mainly include colorized 3D mesh, octree implicit representation, and dense depth supervision, our ablation study analyses the impact of our designs from these aspects. We train five different variations: 1) TCLC-GS without 3D mesh, initializing 3D Gaussian using original 3D LiDAR points; 2) TCLC-GS without colorized 3D mesh, initializing 3D Gaussian using 3D mesh without color information; 3) TCLC-GS without octree implicit representation, initializing 3D Gaussian using colorized 3D mesh without octree representation; 4) TCLC-GS without dense depth supervision, training TCLC-GS only using RGB supervision; 5) TCLC-GS full method. 6) TCLC-GS + sky refine The average values of the evaluation metrics across six testing scenes from the Waymo dataset are given in Table 9. These ablation results indirectly validate the effectiveness of colorized 3D mesh, octree implicit representation, and dense depth supervision, underscoring their contributions to the overall performance.
Running-time Analysis: We present a running time performance comparison between our TCLC-GS and various baselines on the Waymo and nuScenes datasets, detailed in Table 9 and Table 9, respectively. Utilizing a single NVIDIA GeForce RTX 3090 Ti, TCLC-GS attains real-time rendering speeds with high-resolution images, achieving around 90 FPS with an image resolution of 1920$\times$1280 on the Waymo dataset, and around 120 FPS with an image resolution of 1600$\times$900 on the nuScenes dataset. Compared to NeRF-based methods such as SUDS [25] and S-NeRF [12], TCLC-GS significantly outperforms the speed of these methods, achieving real-time running performance. When compared to 3D-GS [11], TCLC-GS offers higher accuracy in both RGB and depth rendering, and still maintains the performance efficiency.
5 Conclusion
In this Appendix, we firstly provide more details of our TCLC-GS implementation, including network architecture, and hyper-parameters of octree feature grid and 3D Gaussians in Appendix 6. Secondly, more comparison visualizations with the baseline are presented in Appendix 7. Lastly, three video demos are attached in Appendix 8.
6 Implementation Details
7 More Visualization Results
Visualization Results on the Waymo Open dataset: We present visual comparison results of image and depth synthesis on the Waymo dataset in Fig. 8 and Fig. 10. It can be seen that TCLC-GS produces clearer images compared to 3D-GS [11]. Additionally, the depth images generated by TCLC-GS are denser and sharper than those produced by 3D-GS.
@@ 8 Video Demo
We provide Video Demo 1, Demo 2 and Demo 3 in the website  [^1] https://github.com/BoschRHI3NA/Video-Demo-ECCV24/tree/main. Demos 1 and 2 illustrate the rendered images and depths by our TCLC-GS, alongside the ground truth (GT) data, in two different urban scenes. In these videos, the first row displays the images rendered by TCLC-GS, the second row shows the GT images, the third row presents the dense depth rendered by TCLC-GS, and the fourth row illustrates the GT sparse depth obtained by projecting LiDAR data onto the surrounding images. Demo 3 showcases a more challenging rendering scenario. This video presents image and depth rendering from a new generated ego-car trajectory which is 0.5 meter higher or 0.5 meter lower than the original ego-car trajectory. The first and third rows display the rendered RGB and depth images from a trajectory 0.5m lower than the original, while the second and fourth rows show the rendered RGB and depth images from a trajectory 0.5m higher than the original. This demonstration highlights our method’s potential to address the domain gap in data reuse caused by different hardware car settings, such as collecting data from a sedan but deploying it on an SUV.